OpenStudy (anonymous):

Solve x2 + 4x + 8 = 0. Solve x2 - 7x = -13 Solve -3x2 - 4x - 4 = 0 Solve 3x2 = -12x - 15 Solve -3x2 + 30x - 90 = 0

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

I will help you with one of them..Then you try the rest. They are all similar. Which one should I help you with?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The first one please! @Easyaspi314

OpenStudy (anonymous):

._. Really?

OpenStudy (anonymous):

x^2 + 4x + 8 = 0 This does NOT factor, so we'll use the Quadratic Formula x = -b +/- sqrt(b^2 - 4ac)/2a In our problem, a = 1, b = 4, c = 8 so x = (-4 +/- sqrt(-16))/2 = -4 +/- 4i)/2 x = -2 +/- 2i are the two roots. NOTE: sqrt (-16) is an imaginary number equal to 4i.